# Nightjar: Dynamic Adaptive Speculative Decoding for Large Language Models Serving

### v2

@@ Abstract
Speculative decoding (SD) accelerates LLM inference by verifying draft tokens in parallel. However, this method presents a critical trade-off: it improves throughput in low-load, memory-bound systems but degrades performance in high-load, compute-bound environments due to verification overhead. Existing Speculative Decoding strategies typically rely on static speculative lengths, failing to adapt to fluctuating request loads or identify the optimal moment to halt speculation. The cost of restarting speculative inference also remains unquantified. During traffic surges, the marginal utility of speculation diminishes; yet, the draft model’s persistent memory footprint competes for available KV cache. This resource contention limits the maximum batch size, thereby degrading overall system throughput. To overcome this, we propose Nightjar, a novel learning-based algorithm for adaptive speculative inference that adjusts to request load by dynamically selecting the optimal speculative length for different batch sizes. Upon detecting significant request queuing and KV cache shortages, we disable speculative decoding and offload the draft model to the CPU. This reclaims VRAM for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Experiments show that Nightjar achieves up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating robust efficiency for real-time serving.
@@ 1 Introduction
Figure 1: Schematic roofline illustration of autoregressive decoding (w/o SD) and speculative decoding (SD); positions are illustrative.
However, current speculative decoding implementations (e.g., in vLLM [13]) often use a fixed speculative length, which is not optimal for dynamic request loads. The optimal speculative length is a dynamic variable dictated by the interplay of batch size, token acceptance rates, and hardware characteristics. An excessively long speculative length increases the time spent on draft generation and imposes significant verification overhead, whereas a length that is too short fails to fully capitalize on the potential speedup of speculative decoding.
Furthermore, model-based draft models consume both compute resources and precious GPU memory (VRAM), unlike model-free approaches such as N-gram that incur negligible memory footprints. However, model-free methods lack generalizability and are often restricted to specific domains like code generation [20]. In high-traffic scenarios, the VRAM occupied by these draft models directly competes with the KV cache; reclaiming this space for the KV cache would facilitate larger batch sizes and significantly maximize overall system throughput.
@@
Therefore, the field urgently needs a truly intelligent, adaptive decision-making mechanism that can perceive load changes in real-time and accurately weigh the trade-offs between benefits and costs. It must be capable of dynamically selecting the optimal speculative strategy for different batch sizes at each decoding step—including decisively disabling it when necessary. To address this core challenge, we propose Nightjar, a contextual bandit approach for speculative length selection in dynamic, real-time LLM services. A key feature of our approach is its ability to learn the optimal speculative length for different batch sizes at each decoding step, without requiring prior knowledge of request difficulty or model characteristics. Furthermore, we incorporate the switching cost of transitioning from a speculative length of 0 to a positive value, a factor previously overlooked. By dynamically offloading the draft model upon disabling speculative inference, we reclaim substantial VRAM for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Simultaneously, when a reduction in request load is detected, we proactively and asynchronously reload the draft model to re-enable speculative decoding, ensuring sustained performance gains. Implemented on vLLM [13] and evaluated on three real-world datasets, Nightjar consistently outperforms existing methods, achieving up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating its practical effectiveness for real-time LLM serving.
@@ 2.2 Memory underutilization
Context Length Vicuna-13B (MHA) DeepSeek-R1-Distill-Qwen-7B
@@
As evidenced by Table I, the memory footprint of the KV cache is not only substantial but scales aggressively with context length, transitioning from a negligible overhead to a primary consumer of GPU resources. This expansion creates a zero-sum competition within the finite VRAM budget: as the KV cache encroaches upon the available memory, it directly constrains the headroom for resident draft model weights. Such resource contention highlights a critical bottleneck in static memory partitioning, where the burgeoning cache demands can eventually marginalize auxiliary models and degrade overall system throughput.
@@
As shown in Figure 3, draft models (Ngram, DeepSeek-R1-DRAFT-Qwen2.5-0.5B [2]) occupy varying GPU memory, impacting KV cache capacity. While their weights are smaller than the main model, they constrain memory under high loads. For instance, unloading a 1.4GB draft model (e.g., 14B Multi-Token Prediction (MTP) for DeepSeek 671B [9]) frees space for  49k additional tokens in KV cache (calculated via MLA [9]). Current systems (e.g., vLLM [13]) retain draft models even when SD is disabled, wasting memory.
@@
We identify that memory inefficiency arises from the isolation of weight and KV cache spaces in existing systems. This isolation hinders the dynamic redistribution of memory resources based on workload demands, leading to underutilized activation memory, which exacerbates KV cache space pressure and ultimately constrains overall system performance.
3 Nightjar Design
In this section, we first introduce the problem statement and the key notations used in Nightjar. Then, we present the algorithm design and the elastic memory management mechanism for Speculative Decoding.
3.1 Overview
Figure 4: An overview of our adaptive approach with multi-armed bandit (MAB) for dynamic request load.
Nightjar consists of two core components: 1) Speculative Length Selection: A contextual multi-armed bandit (MAB) algorithm that selects the optimal speculative length for each batch size in continual batching. If the speculative length is selected to be 0, the draft model will not be used and the system will switch to the standard autoregressive decoding. Otherwise, the draft model will be used to generate the speculative length tokens, and the target model will verify them in parallel. 2) Elastic Memory Management: A mechanism that dynamically allocates memory for the draft model. If the draft model is not used, and request rate is high, and there is not enough GPU memory for KV cache to allocate, we free the draft model in the GPU and make room for the KV cache. When the request rate is low, we load the draft model back into the GPU to accelerate the generation, as illustrated in Figure 4.
@@ 4.1 Problem Statement
In continual batching [13] for LLM serving, the batch size can vary at each decoding step. We consider batch sizes $B\in\mathcal{B}=\{1,2,\dots,B_{\max}\}$ and speculative lengths $\gamma\in\{0,1,\dots,\Gamma_{\max}\}$. Following [19], we evaluate performance using goodput $g_{B,\gamma}$, defined as the number of tokens generated per second, including both accepted proposed tokens and bonus tokens generated during verification.
Our objective is to select, for each batch size $B$, the speculative length $\gamma$ that maximizes goodput:
$\gamma^{*}(B)=\mathop{\arg\max}_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}g_{B,\gamma}$ (1)
However, $g_{B,\gamma}$ depends on the token acceptance rate, which is unknown a priori. We model the problem as a regret minimization task over T steps. We define the reward $r^{t}_{\gamma}$ as the realized goodput observed after selecting arm $\gamma$ at step $t$. The cumulative regret is:
$R(T)=\sum_{t=1}^{T}\left(r^{t}_{\gamma^{*}_{t}}-r^{t}_{\gamma_{t}}\right)$ (2)
@@
$c_{\mathrm{prefill}}$ Latency overhead for reconstructing the KV cache (switching cost).
$L_{\max}$ Effective skip length, defined as $\max_{i\in B}L_{i}$ (max lag in a batch).
$C_{\text{switch}}$ Upper bound constant for the switching cost.
Elastic Memory Management
$M_{\text{orig}},M_{\text{scale}}$ Baseline KV cache capacity and expanded capacity after offloading.
$S_{\text{draft}}$ Memory footprint required to store the draft model weights.
$N_{\text{free}}$ Number of currently available physical KV cache blocks.
$\tau_{\text{low}}$ Critical scarcity threshold triggering memory expansion.
$|\mathcal{Q}_{\text{wait}}|$ Cardinality of the request waiting queue.
@@
$b_{B}$ Bin index within block $j_{B}$, controlling exploration probability ($p\propto 1/\sqrt{b_{B}}$).
@@
$g_{B,\gamma},\tilde{g}_{B,\gamma}$ Theoretical goodput and its empirical mean estimate.
$\gamma^{*}_{t}$ Optimal speculative length at time $t$ that minimizes the objective.
$r_{t}(\gamma)$ Realized reward (observed goodput) at step $t$ for arm $\gamma$.
$N_{t}(\gamma)$ Number of times arm $\gamma$ has been selected up to time $t$.
$\beta_{t}(\gamma)$ Confidence radius for the goodput estimate at time $t$.
$R(T)$ Cumulative regret over horizon $T$ relative to the optimal policy.
$\delta$ Confidence parameter, where estimates hold with prob. $1-\delta$.
@@ 4.2 Algorithm Design
Contextual MAB can take dynamic request batch size or input sequences as context, such as linear Thompson [1], LinUCB [16], which have been successfully applied in online prediction settings such as configuration selection [24]. However, they are questionable in dynamic serving environments. We adopt the relatively simple framework ADA-BINGREEDY proposed in [22]. One difference between [22] and our approach is that we develop a request batch mechanism to handle varying request patterns. And we design reward that considers the speculative length switch cost.
@@
The decision-making process in Nightjar is driven by an objective function that minimizes the inverse goodput while penalizing frequent re-initializations of the draft model. Formally, at global time $t$ with batch size $B$, the optimal speculative length $\gamma_{t}^{*}$ for an exploitation step is determined by:
$\gamma_{t}^{*}=\mathop{\arg\min}\limits_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}\left\{\frac{1}{\tilde{g}_{B,\gamma}}+\frac{\mathbb{I}(\gamma_{t-1}=0\land\gamma>0)\cdot c_{\mathrm{prefill}}}{\gamma}\right\}$ (3)
Here, we estimate the empirical mean goodput $\tilde{g}_{B,\gamma}$ using an incremental cumulative moving average to ensure computational efficiency without storing historical reward sequences. Upon observing a new reward $r_{t}$ for a specific batch size $B$ and speculative length $\gamma$ at the $n$-th visit, the statistic is updated iteratively via $\tilde{g}_{B,\gamma}^{(n)}\leftarrow\tilde{g}_{B,\gamma}^{(n-1)}+\frac{1}{n}(r_{t}-\tilde{g}_{B,\gamma}^{(n-1)})$, where $\tilde{g}_{B,\gamma}^{(0)}=0$.
The first term of Eq. (3), $1/\tilde{g}_{B,\gamma}$, is mathematically dual to maximizing the amortized throughput. The second term functions as a regularization mechanism for the switching cost. The indicator function $\mathbb{I}(\cdot)$ activates only when the system transitions from a disabled state ($\gamma_{t-1}=0$) to an enabled state ($\gamma>0$), incurring a KV cache reconstruction cost $c_{\mathrm{prefill}}$. By dividing this cost by $\gamma$, the objective favors configurations that amortize the startup overhead across longer speculative sequences, thereby discouraging short-sighted switching that degrades system latency.
@@
Initialization and Bin Selection. For every time step $t$, the algorithm identifies the current batch size $B$ and determines the strategy for the current bin. As shown in Lines 1-1, at the onset of a new bin (when the round counter $\tau_{B}=1$), the bin is designated as an exploration bin with probability $1/\sqrt{b_{B}}$ and as an exploitation bin otherwise. This decaying exploration probability ensures that the system explores aggressively in the early stages to gather statistics but progressively converges to the optimal policy as the bin index $b_{B}$ increases.
Execution and State Evolution. During an exploration bin, the speculative length $\gamma_{t}$ is sampled uniformly from the available search space to ensure coverage (Lines 11). Conversely, in an exploitation bin, the algorithm selects $\gamma_{t}$ by solving Eq. (3) to maximize system efficiency (Lines 11). Following the execution of $\gamma_{t}$, the system observes the reward $r_{t}(\gamma_{t})$ and updates the internal counters (Line 1).
@@
Prefill Cost Modeling The effectiveness of Eq. (3) relies on an accurate estimation of $c_{\mathrm{prefill}}$, which represents the KV cache reconstruction latency. This cost is not static; it varies based on the hardware state and the synchronization gap between the draft and target models. When speculative decoding is re-enabled, the draft model must process all tokens generated by the target model during the inactive period.
TABLE III: Measured switching cost ($c_{\text{prefill}}$) across varying batch sizes and input lengths using a 7B model on an NVIDIA RTX 4090.
Input Length Batch Size $c_{prefill}$ (ms)
@@
We quantify this overhead using the effective skip length, defined as $L_{\text{max}}=\max\{L_{1},\dots,L_{B}\}$, where $L_{i}$ is the number of new tokens for request $i$. Since the draft model computes the tokens in parallel, the batch latency is determined by the request with the largest lag. The algorithm queries an offline-populated lookup table $c_{\mathrm{prefill}}(L_{\text{max}},B)$ (Table III) to obtain the precise cost. For example, empirical profiling on a DeepSeek-R1-Distill-Qwen-7B model target model and 0.5B draft model on RTX 4090 indicates that $c_{\mathrm{prefill}}$ can range from 17.87 ms to over 102.03 ms depending on $L_{\text{max}}$, making this dynamic cost retrieval essential for robust decision-making. To construct this lookup table, we performed offline profiling. We measured latencies across a grid of input lengths and geometrically increasing batch sizes (e.g., 2, 4, 8…). Specifically, we recorded the prefill time for the standalone target model ($T_{\text{base}}$) and the prefill time with speculative decoding enabled ($T_{\text{SD}}$). The switching cost $c_{\text{prefill}}$ is derived from the difference ($T_{\text{SD}}-T_{\text{base}}$).

### v3

@@ Abstract
Speculative decoding (SD) accelerates LLM inference by verifying draft tokens in parallel. However, this method presents a critical trade-off: it improves throughput in low-load, memory-bound systems but degrades performance in high-load, compute-bound environments due to verification overhead. Existing Speculative Decoding strategies typically rely on static speculative lengths, failing to adapt to fluctuating request loads or identify the optimal moment to halt speculation. The cost of restarting speculative inference also remains unquantified. During traffic surges, the marginal utility of speculation diminishes; yet, the draft model’s persistent memory footprint competes for available KV cache. This resource contention limits the maximum batch size, thereby degrading overall system throughput. To overcome this, we propose Nightjar, a resource-aware adaptive speculative framework. It first adjusts to the request load by dynamically selecting the optimal speculative length for different batch sizes. Crucially, upon detecting significant request queuing or KV cache shortages, it disables speculative decoding and offloads the draft model to the CPU. This reclaims memory for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Experiments show that Nightjar achieves up to 27.29% higher throughput and 12.90% lower latency compared to standard speculative decoding under dynamic request arrival rates in real-time LLM serving scenarios.
@@ 1 Introduction
Figure 1: Conceptual Roofline model illustrating the performance shift in LLM inference between standard autoregressive decoding (w/o SD) and speculative decoding (SD); positions are illustrative.
However, current speculative decoding implementations (e.g., in vLLM [14]) often use a fixed speculative length, which is not optimal for dynamic request loads. The optimal speculative length is a dynamic variable dictated by the interplay of batch size, token acceptance rates, and hardware characteristics. An excessively long speculative length increases the time spent on draft generation and imposes significant verification overhead, whereas a length that is too short fails to fully capitalize on the potential speedup of speculative decoding. Crucially, an effective system must decisively disable speculation when its marginal benefits vanish.
Furthermore, model-based draft models consume both compute resources and precious GPU memory, unlike model-free approaches such as N-gram that incur negligible memory footprints. However, model-free methods lack generalizability and are often restricted to specific domains like code generation [22]. In high-traffic scenarios, the memory occupied by these draft models directly competes with the KV cache; reclaiming this space for the KV cache would facilitate larger batch sizes and significantly maximize overall system throughput.
@@
Therefore, the field urgently needs a truly intelligent, adaptive decision-making mechanism that can perceive load changes in real-time and accurately weigh the trade-offs between benefits and costs. It must be capable of dynamically selecting the optimal speculative strategy for different batch sizes at each decoding step, including decisively disabling it when necessary. To address this core challenge, we propose Nightjar, a contextual bandit approach for speculative length selection in dynamic, real-time LLM services. A key feature of our approach is its ability to learn the optimal speculative length for different batch sizes at each decoding step, without requiring prior knowledge of request difficulty or model characteristics. Furthermore, we incorporate the switching cost of transitioning from a speculative length of 0 to a positive value, a factor previously overlooked. By dynamically offloading the draft model upon disabling speculative inference, we reclaim substantial memory for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Simultaneously, when a reduction in request load is detected, we proactively and asynchronously reload the draft model to re-enable speculative decoding, ensuring sustained performance gains. Implemented on vLLM [14] and evaluated on three real-world datasets, Nightjar significantly accelerates LLM inference, establishing a new state-of-the-art. It consistently outperforms existing methods, achieving an average throughput improvement of 27.29% over the vanilla baseline and 7.39% over the strongest existing method. Coupled with up to 12.90% lower latency compared to standard speculative decoding, these results demonstrate Nightjar’s practical effectiveness for dynamic and real-time LLM serving.
@@ 2.2 Memory underutilization
Context Length Vicuna-13B DeepSeek-R1-Distill-Qwen-7B
@@
As evidenced by Table I, the memory footprint of the KV cache is not only substantial but scales aggressively with context length, transitioning from a negligible overhead to a primary consumer of GPU resources. This expansion creates a zero-sum competition within the finite memory budget: as the KV cache encroaches upon the available memory, it directly constrains the headroom for resident draft model weights. Such resource contention highlights a critical bottleneck in static memory partitioning, where the burgeoning cache demands can eventually marginalize auxiliary models and degrade overall system throughput.
@@
As shown in Figure 3, draft models (Ngram, DeepSeek-R1-DRAFT-Qwen2.5-0.5B [2], Eagle [19]) occupy varying GPU memory, impacting KV cache capacity. While their weights are smaller than the main model, they constrain memory under high loads. This memory overhead is particularly critical for large-scale models. Consider a standard deployment of DeepSeek-V3 671B [10] across $\timesH10$0 (80GB) GPUs. The 14B MTP draft model, distributed via tensor parallelism, consumes 1.4GB of HBM per GPU. Under the MLA architecture [10], this specific memory slice is extremely valuable, equivalent to the storage of $\sim$49k KV cache tokens. Current systems (e.g., vLLM [14]) retain draft models even when SD is disabled, wasting memory.
@@
Fundamentally, we identify that this memory inefficiency arises from the isolation of weight and KV cache spaces in existing systems. This isolation hinders the dynamic redistribution of memory resources based on workload demands, leading to underutilized activation memory, which exacerbates KV cache space pressure and ultimately constrains overall system performance.
3 System Architecture
Figure 4: The Nightjar architecture operates across three layers: (1) The Scheduler manages incoming requests; (2) The Planner makes step-wise decisions to toggle speculation; (3) The Memory Manager dynamically reallocates resources between model weights and the KV cache pool based on system load.
As illustrated in Figure 4, Nightjar is designed as a vertically integrated framework that coordinates decision-making across three functional layers: the Scheduler, Engine Execution, and Memory Management.
1) Scheduler and Planner Layer. At the top level, the Request Scheduler handles continuous batching for incoming requests. It feeds real-time batch statistics (e.g., current batch size $B$) to the Nightjar Planner, which employs a Contextual Multi-Armed Bandit (MAB) engine to determine the optimal speculative length $\gamma$ for the current decoding step.
2) Engine Execution Layer. Driven by the Planner’s decision, the execution engine dynamically switches between two modes:
  • Speculative Decoding (Right): When the Planner selects $\gamma>0$, the system engages the draft model to propose tokens, which are verified in parallel by the target model.
  • Standard AR Decoding (Left): If the Planner selects $\gamma=0$, speculation is disabled. The system bypasses the draft model entirely to avoid verification overhead.
3) Memory Management Layer. Underlying the execution is an elastic Memory Manager. As shown in the bottom panel of Figure 4, this module manages the contention between draft model weights and the KV cache. It implements a ”Squeeze/Expand” mechanism: when the speculative decoding is disabled and memory is insufficient for the KV cache, it offloads the draft model to host memory to expand the KV cache pool, maximizing batch size. Conversely, during low load, it reallocates memory to reload the draft model before re-enabling speculative decoding, ensuring accelerated inference.
This hierarchical design allows Nightjar to adapt to fluctuating workloads at both the algorithmic level and the system level.
@@ 4.1 Problem Statement
In continuous batching [14] for LLM serving, the batch size can vary at each decoding step. We consider batch sizes $B\in\mathcal{B}=\{1,2,\dots,B_{\max}\}$ and speculative lengths $\gamma\in\{0,1,\dots,\Gamma_{\max}\}$. While performance is often evaluated using goodput $g_{B,\gamma}$ (tokens/sec), directly maximizing goodput ignores the heavy latency overhead caused by KV cache reconstruction during strategy switches.
To explicitly account for this cost, we formulate the problem as minimizing the latency per token, denoted as $\ell_{B,\gamma}=1/g_{B,\gamma}$. Our objective is to select, for each batch size $B$, the speculative length $\gamma$ that minimizes the effective latency:
$\gamma^{*}(B)=\mathop{\arg\min}_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}\ell_{B,\gamma}$ (1)
Since the token acceptance rate is unknown a priori, we model this as a regret minimization task over $T$ steps. We define the Loss observed at step $t$, denoted as $L_{t}(\gamma)$, to be the realized latency per token plus the switching overhead:
$L_{t}(\gamma_{t})=\ell^{t}_{B,\gamma_{t}}+\mathbb{I}(\gamma_{t-1}=0\land\gamma_{t}>0)\cdot\frac{C_{\text{switch}}}{\gamma_{t}}$ (2)
where $\ell^{t}_{B,\gamma_{t}}$ is the observed inverse goodput, $\mathbb{I}(\cdot)$ is the indicator function for a strategy switch, and $C_{\text{switch}}$ is the latency overhead of KV cache reconstruction amortized over the generated tokens.
The cumulative regret $R(T)$ is defined as the difference between the total loss incurred by the algorithm and the loss of the optimal fixed policy $\pi^{*}$ that knows the true distributions:
$R(T)=\sum_{t=1}^{T}\left(L_{t}(\gamma_{t})-L_{t}(\gamma^{*})\right)$ (3)
@@
$L_{\max}$ Effective skip length, defined as $\max_{1\leq i\leq B}L_{i}$ (max lag in a batch).
$C_{\text{switch}}$ Constant representing the KV cache reconstruction cost in the loss function.
@@
$b_{B}$ Bin index within block $j_{B}$, controlling exploration probability ($p\propto 1/b_{B}$).
@@
$g_{B,\gamma},\tilde{g}_{B,\gamma}$ Theoretical goodput (tokens/s) and its empirical mean estimate.
$\ell_{B,\gamma}$ Expected latency per token, defined as the inverse goodput ($1/g_{B,\gamma}$).
$\gamma^{*}$ Optimal speculative length that minimizes the total loss.
$L_{t}(\gamma)$ Realized loss at step $t$, including latency and switching penalty.
$R(T)$ Cumulative regret (in terms of latency) over horizon $T$.
Elastic Memory Management
$M_{\text{orig}},M_{\text{scale}}$ Baseline and expanded KV cache memory size (bytes); $M_{\text{scale}}=M_{\text{orig}}+S_{\text{draft}}$.
$S_{\text{draft}}$ Draft model weight memory size (bytes).
$B_{\text{block}}$ Bytes per KV cache block.
$N_{\text{orig}},N_{\text{scale}},N_{\text{draft}}$ KV cache capacity in blocks: baseline, expanded, and draft-equivalent ($N_{\text{draft}}=\lceil S_{\text{draft}}/B_{\text{block}}\rceil$).
$N_{\text{free}}$ Number of currently available physical KV cache blocks.
$\tau_{\text{low}}$ Critical scarcity threshold (in blocks) triggering memory expansion.
$|\mathcal{Q}_{\text{wait}}|$ Cardinality of the request waiting queue.
@@ 4.2 Algorithm Design
Contextual MAB can take dynamic request batch size or input sequences as context, such as linear Thompson [1] or LinUCB [17], which have been successfully applied in online prediction settings like configuration selection [26]. However, their complexity makes them questionable for latency-sensitive dynamic serving environments. We adopt the relatively simple framework ADA-BINGREEDY proposed in [24]. A key difference between [24] and our approach is that we develop a request batch mechanism to handle varying request patterns, and we design a loss function that explicitly considers the speculative length switching cost.
@@
To explicitly account for the switching overhead, we reformulate the maximization of goodput into the minimization of latency per token. Formally, at global time $t$ with batch size $B$, the optimal speculative length $\gamma_{t}^{*}$ for an exploitation step is determined by minimizing the regularized latency:
$\gamma_{t}^{*}=\mathop{\arg\min}\limits_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}\left\{\frac{1}{\tilde{g}_{B,\gamma}}+\frac{\mathbb{I}(\gamma_{t-1}=0\land\gamma>0)\cdot C_{\text{switch}}}{\gamma}\right\}$ (4)
Here, we directly estimate the empirical mean latency per token $\tilde{\ell}_{B,\gamma}$ to avoid bias from Jensen’s inequality ($\mathbb{E}[1/g]\geq 1/\mathbb{E}[g]$). Upon observing the instantaneous latency $\ell_{t}=1/g_{t}$, the statistic is updated iteratively via $\tilde{\ell}_{B,\gamma}^{(n)}\leftarrow\tilde{\ell}_{B,\gamma}^{(n-1)}+\frac{1}{n}(\ell_{t}-\tilde{\ell}_{B,\gamma}^{(n-1)})$.
The second term functions as a regularization mechanism for the switching cost. $\gamma_{t-1}$ is the speculative length at the previous time step. The indicator function $\mathbb{I}(\cdot)$ activates only when the system transitions from a disabled state ($\gamma_{t-1}=0$) to an enabled state ($\gamma>0$), incurring a KV cache reconstruction cost $C_{\text{switch}}$. By dividing this cost by $\gamma$, the objective favors configurations that amortize the startup overhead across longer speculative sequences, thereby discouraging short-sighted switching that degrades system latency.
@@
Initialization and Bin Selection. For every time step $t$, the algorithm identifies the current batch size $B$. Critical decision-making occurs at the onset of a new bin (when the round counter $\tau_{B}=1$). At this step, the algorithm determines the bin type: it is designated as an exploration bin with probability $1/b_{B}$ and as an exploitation bin otherwise. This decaying exploration probability ensures aggressive exploration in early stages and progressive convergence to the optimal policy as $b_{B}$ increases.
Arm Selection and Execution. Upon initializing the bin ($\tau_{B}=1$), the speculative length $\gamma$ is selected and locked for the entire duration of the bin. If it is an exploration bin, $\gamma$ is sampled uniformly from the available search space to ensure coverage (Lines 11). Conversely, in an exploitation bin, the algorithm selects the optimal $\gamma$ by solving Eq. (4) (Lines 11). For all subsequent steps within the bin, the system executes this fixed $\gamma$, observes the realized goodput $g_{t}$, and updates the internal counters. This locking mechanism is essential for bounding the switching cost.
@@
Prefill Cost Modeling. The effectiveness of Eq. (4) relies on an accurate estimation of $C_{\text{switch}}$, which represents the KV cache reconstruction latency. This cost is not static; it varies based on the hardware state and the synchronization gap between the draft and target models.
TABLE III: Measured switching cost ($C_{\text{switch}}$) across varying batch sizes and input lengths using a 7B model on an NVIDIA RTX 4090.
Input Length Batch Size $C_{\text{switch}}$ (ms)
@@
We quantify this overhead using the effective skip length, defined as $L_{\text{max}}=\max\{L_{1},\dots,L_{B}\}$, where $L_{i}$ is the number of new tokens for request $i$. The algorithm queries an offline-populated lookup table $C_{\text{switch}}(L_{\text{max}},B)$ (Table III) to obtain the precise cost. To construct this table, we performed offline profiling across a grid of input lengths and geometrically increasing batch sizes (e.g., 2, 4, 8, …). Specifically, we recorded the prefill time for the standalone target model ($T_{\text{base}}$) and the prefill time with speculative decoding enabled ($T_{\text{SD}}$). The switching cost is derived from the difference $C_{\text{switch}}=T_{\text{SD}}-T_{\text{base}}$. As shown in Table III, on an RTX 4090 setup, this cost can range from 17.87 ms to over 102.03 ms, making dynamic retrieval essential.